OpenStudy (anonymous):

Two trains are headed toward each other on the same straight track, one at 65mph and the other at 45mph. At time t=0 hours, when the trains are 75 miles apart, an insect flies from one train to the other at a constant 35mph. Upon arriving at the other train, the insect reverses its path and travels to the other train at the constant 35mph in the other directions, repeating the process. At some time t[subscript:c], the trains collide. If, for t is greater than or equal to zero, d(t) miles is the total distance traveled by the insect then the limit as t approaches t[subscript:c] from the left of

OpenStudy (anonymous):

d(t)= ______ miles? (Sorry, it was a long question that I have no clue what to do anything with.

OpenStudy (anonymous):

Solve the following equation for t, the time it takes the trains to collide. 65 t + 45 t = 75 t = 15/22 hrs r t = d The insect travels 35*(15/22) or 23.8636 miles.